3) A) Explain the Fisher Effect b) The current inflation expectation is low at about 1%, if the real rate of interest long term is 2%, what will be the yield on treasury bills based on Fisher Effect?
Answer #1

Nominal Yield = [(1 + Real Rate) * (1 + Inflation Rate)] - 1

= (1 + 0.01) * (1 + 0.02) = 1.0302 - 1 = 0.0302, or 3.02%